Operations Report – May 20th

Crew 195 Operations Report 20.05.18

SOL: 0

Name of person filing report: Chris Davis

Non-nominal systems: Continuing run Diesel generator at night. No service rep for Solar yet.

Notes on non-nominal systems:

Generator (hours run): Overnight approximately 10 plus 4 hours today per Shannon

Solar— SOC 52%

Diesel – Over 50%

Propane – 51 PSI

ATV Fuel (Ethanol free gasoline) – 5 gallons.

Water (trailer) – 500 gallons.

Water (static) – Approximately 250

Trailer to Static Pump used – No

Water (loft) – Static to Loft Pump used – Yes

Water Meter: 1351130

Toilet tank emptied: No
Date last emptied: Previous Crew Sol 5

ATV’s (Honda, 300, 350.1, 350.2, 350.3): all except Honda and 300

ATV Oil Checked?: No

ATV Oil Added?: No

ATV Fuel Used: Approximately 3 gallons

# Hours the ATVs were Used today: 1 hours

Notes on ATVs: Honda and ATV 300 out of service.

Deimos rover used: yes
Hours: N/A
Beginning charge: 96%
Ending charge: 82 %
Currently charging: Yes

Sojourner rover used: ASSIGNED TO DIRECTOR
Hours: Unknown
Beginning charge: N/A
Ending charge: N/A
Currently charging: N/A

Spirit rover used: Yes

Hours: 35.8
Beginning charge: 100 %
Ending charge: 90%
Currently charging: Yes

Opportunity rover used: Yes
Hours: 34.2
Beginning charge: 100 %
Ending charge: 92 %
Currently charging: Yes

Curiosity rover used: No
Hours: 29.8
Beginning charge: 100 %
Ending charge: 90%
Currently charging: Yes

Hab Car used?: No
Where was Hab Car used?:
Why was Hab Car used?:

General notes and comments:

Summary of internet: Working well

Summary of suits and radios: working well

Summary of Hab operations: No problems

Summary of GreenHab operations: N/A

Summary of ScienceDome operations: N/A

Summary of RAM operations: N/A

Summary of health and safety issues: none noted

Questions, concerns and requests to Mission Support: None.